OpenStudy (tanya123):

How to factor this and make it into a hyperbola 9y^2-4x^2-54y+32x-19=0

OpenStudy (tanya123):

ive done this part, 9y^2-4x^2-54y+32x=19

Nnesha (nnesha):

\[\large\rm \frac{ (x-h)^2 }{ a^2} - \frac{ (y-k)^2 }{ b^2} =1\] hyperbola equation first of all you have to write both x term in one parentheses and y values into another parentheses

OpenStudy (tanya123):

9(y^2 idk how to complete this.....

OpenStudy (tanya123):

9(y^2-54y) -(32x) ?

Nnesha (nnesha):

 like this